OpenStudy (cometailcane):

A 525 mL container of chlorine gas has a temperature of 75.9 degrees C. The mass of this sample of gas is 14.4 grams. a. determine the number of moles of chlorine gas from its mass. b. convert your volume to liters and temperature of Kelvin c. using the ideal gas law, determine the pressure exerted from the gas.

OpenStudy (cuanchi):

a) the number of moles can be calculated with the mass (14.4g) and the molecular mass of the Cl2 (35.45 x 2) an the formula "n=m/MM" b) to convert volume from mL to L you can use the conversion factor "1 L = 1000mL" to convert the temperature from C to K use the formula "K=C+273.15". Olso can be googled :) c) the Ideal gas law said \[P \times V = n \times R \times T\] then \[P= \frac{n \times R \times T }{ V }\] were n = number of moles calculated in item a R= the gas constant = 0.082 atm L /mol K T= the temperature in Kelvin converted in item b V= volume in liters converted in item b